• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V.Nelyubin

    Log entry time 11:56:18 on April18,2004

    Entry number 118488

    keyword=Target alarm limit changed

    According J.P.Chen call:

    For Loop2
    (L2_CT_loop_R)
    HiHi alarm limit is changed
    from 87.5 to 100.0 and
    High limit from 87.0 to 95.0.
    L2_T1_Aft_Heater_R
    HiHi alarm limit from 81.0 to 85.
    High alarm limit from 81.0 to 95.



    FIGURE 1